### v1

@@ 2.1 BD Designs in Practice
We searched the academic literature in multiple social sciences to learn how applied researchers implemented the BD design. We conducted our search in four general queries. In our initial Economics query, we included ten leading journals in Economics—Journal of Political Economy, Review of Economic Studies, Econometrica, American Economic Review, Quarterly Journal of Economics, Review of Economics and Statistics, American Economic Journal: Applied Economics, American Economic Journal: Economic Policy, American Economic Journal: Macroeconomics, and American Economic Journal: Microeconomics. In our initial Political Science query, we searched six leading Political Science journals—American Political Science Review, American Journal of Political Science, Journal of Politics, Political Science Research and Methods, British Journal of Political Science, and Quarterly Journal of Political Science. In both cases we searched for terms like “boundary discontinuity”, “multivariate regression discontinuity”, and “multiscore regression discontinuity”, and focused our attention on papers published since $2015$. We also expanded our search to include other articles that were repeatedly cited by the articles collected in our initial queries.
@@
We also searched in the field of Education, as methodological researchers in that area were among the earliest scholars who discussed the multi-score RD design in non-geographic settings. For this search, we started with the seminal methodological papers in Education by Reardon-Robinson_2012_JREE and Wong-Steiner-Cook_2013_JEBS, and collected papers that both cited this methodological papers and were cited by them. Finally, we also included some papers in environmental science, to illustrate uses of the design at the intersection between the natural and social sciences. Our review is not meant to provide an exhaustive list of the literature. This literature is vast, and cannot be fully captured by our approach since we excluded several journals and restricted the time period. Our goal is simply to provide an overview of recent empirical work using the BD design across the social sciences to gauge the most common strategies used for empirical analysis.
@@
The results of our search are summarized in Table 5. Our search yielded 82 published papers. Of these, we found that the overwhelming majority (76 articles or approximately 93%) used a pooling approach, focusing on the average effect along the entire boundary and using the univariate closest distance to the boundary as the running variable. Fewer than 10 articles were explicit in accounting for the heterogeneity along the boundary; of these, only 5 reported effects for different points or segments along the boundary (gonzalez2021cell; grout2011land; Keele-Titiunik_2015_PA; snider2015barriers; velez2019tuning). In sum, our review of the practical literature indicates that most applied researchers adopt a pooling approach, performing localization by grouping all observations sufficiently “close” to the boundary and then estimating a single average treatment effect. These common practices motivate the taxonomy we used in the rest of this chapter.
@@ 3 Pooling-Based Methods
Given a sample of outcomes and bivariate scores, $(Y_{1},\mathbf{X}_{1}),\ldots,(Y_{n},\mathbf{X}_{n})$, a distance function $\mathcal{d}(\cdot,\cdot)$ is used to measure the closeness between any two points in $\mathcal{X}$. The most common example is the Euclidean distance $\mathcal{d}(\mathbf{x}_{1},\mathbf{x}_{2})=\|\mathbf{x}_{1}-\mathbf{x}_{2}\|=\sqrt{(x_{11}-x_{21})^{2}+(x_{12}-x_{22})^{2}}$ for $\mathbf{x}_{j}=(x_{j1},x_{j2})$, $j=1,2$, although in some applications this is not the most appropriate measure. For example, ambrus2020loss are interested in how close residences in a city are to each other in terms of how long it takes for a person to walk between them; this requires calculating a walking distance, which is different from the Euclidean distance which would not take into account, say, that the path must follow city blocks and cannot go through buildings. More broadly, in geographic settings it is common practice to employ specific geodetic distances (Banerjee_2005_Biometrics).
Once a distance function $\mathcal{d}(\mathbf{X}_{i},\mathbf{x})$ has been chosen, researchers who employ the pooling approach define the closest signed distance to the boundary,
@@
for $i=1,\ldots,n$. For a unit $i$, $D_{i}$ measures how far $i$’s location $\mathbf{X}_{i}$ is to the boundary point that is closest to $\mathbf{X}_{i}$, regardless of where that boundary point is located. Following the standard RD design logic, but now expanded to the BD design (a bivariate setting), localization occurs simultaneously along the entire assignment boundary because the univariate distance $D_{i}$ to $\mathcal{B}$ is used to construct the region such that $|D_{i}|\leq h$, which covers the boundary $\mathcal{B}$, where $h$ denotes a bandwidth parameter. This region, called a “tubular neighborhood” in mathematics (federer2014geometric), is a natural generalization of the standard localization approach used in univariate RD designs, where the region is simply $|X_{i}-c|\leq h$ when the score $X_{i}$ is scalar and $c$ is the univariate cutoff.
Figure 2 illustrates the idea graphically. In both cases, the tubular neighborhood represents the area that includes the observations used in the pooling-based analysis. Figure 2(a) shows a “large” bandwidth $h_{1}$, while Figure 2(b) shows a “small” bandwidth $h_{2}$. All units contribute simultaneously, regardless of their specific location. In particular, two units $i$ and $j$ with the same signed distance to the boundary, $D_{i}=D_{j}=d$, will be assigned to the same group (treatment or control, depending on the sign of $d$) and will be $|d|$ units away from the boundary, but could be arbitrarily far from each other.
@@
Figure 3 presents the two resulting RD plots (Calonico-Cattaneo-Titiunik_2015_JASA) in the pooling approach, based on the outcome data $Y_{i}$ and the univariate distance score $D_{i}$ (the cutoff is $c=0$ by construction of $D_{i}$). These plots include a fourth-order global polynomial fit (solid line) for visualization purposes only; those global approximations are not recommended for estimation of treatment effects because they tend to exhibit poor performance due to the Runge’s phenomenon. Instead, as anticipated in Figure 2 and demonstrated in Figure 3, the idea is to localize around the “cutoff” (i.e., the one-dimensional boundary curve) to conduct estimation and inference.
@@
In the pooling approach, the analysis is conducted using all the observations with score $\mathbf{X}_{i}$ within the tubular neighborhood determined by the bandwidth $h$, that is, for all units with $|D_{i}|\leq h$. However, for this subsample, the specific treatment effect estimation approach can vary substantially. To describe and unify the most common approaches in the empirical literature illustrated in Table 5 as well as other approaches recommended from a methodological perspective, we define the following generic weighted least squares regression notation:
@@
where $Y_{i}$ is the outcome variable, $\mathbf{Z}_{i}$ is a (column) vector of independent variables, and $W_{i}$ are weights. A localized regression analysis can be implemented by setting $W_{i}=\mathds{1}(|D_{i}|\leq h)$ or, more generally, $W_{i}=K(D_{i}/h)$ for a kernel function $K(\cdot)$ determining the relative weight given to each observation. The bandwidth $h$ determines the degree of localization, as illustrated in Figure 2. When $\mathbf{Z}_{i}$ contains polynomial expansions of $D_{i}$, the resulting estimation approach is often called nonparametric local polynomial regression but, as we will see shortly, empirical researchers often employ other variables in their local regression fit within the tubular determined by $D_{i}$.
@@
All empirical approaches naturally include $T_{i}$ in $\mathbf{Z}_{i}$ to estimate treatment effects; the majority also include other variables such as $D_{i}$, $\mathbf{X}_{i}$, boundary-segment fixed effects (formally defined below), as well as certain transformations and interactions thereof. For simplicity, we omit discussing the inclusion of predetermined covariates, but we note that researchers often further augment the basic local regression specifications with pretreatment variables such as census characteristics, terrain ruggedness and elevation, time fixed effects, among other features. For a discussion of the role of preintervention covariates in the standard RD design see Calonico-Cattaneo-Farrell-Titiunik_2019_RESTAT for efficiency gains and Calonico-Cattaneo-Farrell-Palomba-Titiunik_2025_wp for treatment effect heterogeneity.
The pooling approach for analyzing BD designs begins by choosing the localization bandwidth $h$ to determine the units with $|D_{i}|\leq h$ that will be used in the subsequent estimation. Until now, there were no formal methods for choosing the bandwidth $h$ in an objective and data-driven way, so researchers often considered a range of “reasonable” values. For example, black1999better reported results in three different bandwidths, keeping observations within $0.35$, $0.25$, and $0.15$ miles of the nearest point on the boundary. As in standard univariate RD designs, different choices of $h$ trade off bias and variance in the estimation, with a larger $h$ leading to more bias but less variance, and a smaller $h$ having the opposite effect. Cattaneo-Titiunik-Yu_2025_BDD-Pooling recently established formal results characterizing this tradeoff and provided formal guidance for choosing the bandwidth when implementing pooling BD methods. We discuss these methods in Section 3.1.
Once the bandwidth $h$ has been chosen, researchers pool all units with distance $D_{i}$ within $h$ and then compare those assigned to the treatment group ($D_{i}\geq 0$) to those assigned to the control group ($D_{i}<0$). Because the bandwidth localizes the analysis to only a “small” region on either side of the boundary $\mathcal{B}$, the empirical analysis compares barely treated units to barely control units which, as in the standard univariate RD design, is the basis for the causal interpretation of the comparison, as all confounders are assumed to vary smoothly at the boundary while the treatment assignment changes abruptly from zero to one.
The most basic analysis proceeds by assuming that the units within the shrinking tubular of “width” $2h$ are as-if randomly assigned to treatment and control. This approach is akin to the “local randomization” framework in standard RD designs (Lee-Lemieux_2010_JEL; Cattaneo-Frandsen-Titiunik_2015_JCI; Cattaneo-Titiunik-VazquezBare_2017_JPAM). Thus, a natural treatment effect estimator is the difference-in-means of the outcome between treated and control units with distance to the boundary no larger than the bandwidth:
@@
For example, black1999better reported results using specification (1), calling it a “rough, nonparametric estimate of the value of better schools” (p. 590). She also augmented the basic specification to include “segment fixed effects”, that is, binary variables indicating whether an observation is closest to one of several segments partitioning the assignment boundary. To describe this approach formally, we let $\mathcal{B}_{\ell}$, $\ell=1,\ldots,L$, be disjoint segments that partition the assignment boundary, that is, $\mathcal{B}=\sqcup_{1\leq\ell\leq L}\mathcal{B}_{\ell}$. Then, $S_{i}=\operatorname*{arg\,min}_{1\leq\ell\leq L}d(\mathbf{X}_{i},\mathcal{B}_{\ell})$ is the boundary-segment allocation random index for each unit $i=1,\ldots,n$, where it is understood that when the $\operatorname*{arg\,min}$ contains more than one element, the smallest item is chosen. The vector of segment fixed effects for observation $i$ is $\boldsymbol{\iota}_{L}(S_{i})=(\mathds{1}(S_{i}=1),\cdots,\mathds{1}(S_{i}=L))$, which collects the $L$ boundary segment binary indicators determined by $i$’s location. The difference-in-means specification with segment fixed effects is
@@
where the intercept is removed due to its perfect collinearity with the boundary-segment fixed effects. Setting $L=1$ reduces $\boldsymbol{\iota}_{L}(S_{i})$ to the common intercept in specification (1).
@@
For example, black1999better used a modified version of specification (2) where she included the test scores of the school associated with observation $i$ instead of the binary treatment assignment indicator $T_{i}$. The causal interpretation of this specification rests on the assumption that, after localizing the analysis to include only observations near the boundary of school attendance districts, the only factor that can explain a difference in average house prices between the treated and control areas is school quality, because school quality changes discontinuously at the boundary while all other determinants of house prices are equal on average. In this sense, using test scores instead of the treatment assignment indicator in specifications (1) and (2) is in the spirit of an instrumental variables strategy where $T_{i}$ is used as an instrument for test scores (which proxy school quality), and the localization to a small bandwidth around the boundary ensures that units assigned to treatment and units assigned to control are similar in all other characteristics that might affect house prices.
@@
The seminal research design used by black1999better is prior to the development of modern methods for RD analysis, which occurred predominantly during the $2000$s; see Cattaneo-Titiunik_2022_ARE for an overview. In more recent practice, the simple specifications (1) and (2) are not commonly used because researchers tend to adopt a continuity-based approach rather than a local randomization approach for interpretation. Under the continuity-based approach, the local constant fit has poor control of the misspecification bias, and hence practitioners prefer more flexible local regression specifications.
@@
Later empirical work using BD designs made more explicit the connection between geographic discontinuities in treatment assignment and the classical RD design, following more closely the methods of univariate continuity-based RD analysis, which emphasized including the RD score in the polynomial specification to allow for arbitrary dependence between the RD score and outcome even within the bandwidth. Although there is considerable diversity in the specifications used in more recent practice, most researchers that include the score adopted one of two main strategies: a polynomial fit of the outcome on the bivariate score $\mathbf{X}_{i}$, or a polynomial fit of the outcome on the univariate distance $D_{i}$.
@@
In an influential contribution, dell2010persistent augmented specification (2) by including the bivariate score $\mathbf{X}_{i}$. She studied the long-run impact of the mita, a forced mining labor system instituted by the Spanish in Peru and Bolivia between $1573$ and $1812$, using the historical boundary that determined which communities were forced to send labor to define treated and control areas. She studied the effect of the mita on contemporary outcomes measured at the household level, including consumption and childhood stunting. Her main local regression specification was
@@
which includes the treatment indicator, boundary segment fixed effects, and a polynomial expansion of the bivariate score $\mathbf{X}_{i}$ (with $p=3$ in her case). The estimated coefficient of interest is always the coefficient that accompanies the treatment indicator $T_{i}$. In our notation, $\mathbf{r}_{p}(\mathbf{u})=(u_{1},u_{2},u_{1}^{2},u_{1}u_{2},u_{2}^{2},\ldots,u_{1}^{p},u_{1}^{p-1}u_{2},\ldots,u_{1}u_{2}^{p-1},u_{2}^{p})$ with $\mathbf{u}=(u_{1},u_{2})$. The inclusion of (a polynomial basis of) the bivariate location score $\mathbf{X}_{i}$ is meant to control for “the smooth effects of geographic location” dell2010persistent. A notable characteristic of this specification is that it does not include interactions between the bivariate score and the treatment indicator, which imposes the restriction that the polynomials be identical in the treated and control areas. As for bandwidth selection, dell2010persistent considered 100100km of the mita boundary first, but this choice is later reduced to 7575km and 5050km in subsequent analyses.
@@
Specification (3) has become a common strategy to analyze BD designs. For example, mendez2022multinationals use it to study the effect of private sector companies in the development of local amenities, using a land concession to the multinational United Fruit Company in Costa Rica; paulsen2023foundations use it to estimate the effect of school funding on future earnings and political participation in New York, comparing areas that received higher funding to neighboring areas that did not; and de2018agents compare areas in South Africa where traditional leaders remain highly influential to areas where they are not, and estimate the electoral impact of strong political ties between traditional leaders and the African National Congress (ANC). Other examples are listed in Table 5. All of these examples follow specification (2) by including a flexible polynomial expansion of latitude and longitude, as well as boundary-segment fixed effects, but without allowing for interactions with the treatment indicator $T_{i}$.
@@
with the polynomial expansion reducing to the simpler vector of the univariate distance $\mathbf{r}_{p}(D_{i})=(D_{i},D_{i}^{2},\ldots,D_{i}^{p})$. As in the case of specification (3), specification (4) can also be augmented with boundary-segment fixed effects:
@@
where the estimated coefficient of interest continues to be the one accompanying $T_{i}$. In practice, specification (6) is often implemented with $p=1$ and, depending on whether standard software for RD design analysis is used, the specification may employ a non-uniform weighting kernel (e.g., a triangular kernel that down-weights observations as their distance from the boundary increases). Putting aside the inclusion of boundary-segment fixed effects, Figure 6 illustrates the importance of including the interaction term $T_{i}\cdot\mathbf{r}_{p}(D_{i})$ in the treatment effect estimation, as in specification (6). Comparing Figures 5 and 6 shows that the two approaches can lead to different local approximations, with specifications (1) through (5) potentially exhibiting a larger bias than specification (6).
@@
Specification (6) is one of the most commonly used in recent practice. Examples include ayres2021environmental, eugster2017culture, jones2022factor, lalive2008extended, and lowes2021concessions. We note that several articles write their empirical specifications as (4) or (5), but they actually implement specification (6) in their analyses (as clarified in text, footnotes or replication files). These researchers employ standard practices and software from the univariate RD design literature, which follows specification (6).
@@
where the estimated treatment effect coefficient accompanying $T_{i}$ enjoys better bias properties due to the added flexibility in the local regression specification. However, this more flexible approach is not as common in the empirical literature—some exceptions are gonzalez2021cell, egger2015impact, and jones2022effects.
In some applications, researchers estimate treatment effects in BD designs by considering two separate univariate RD designs, that is, keeping one dimension fixed and studying the other dimension of the bivariate score. For example, LondonoVelezRodriguezSanchez_2020_AEJ investigate the effects of Ser Pilo Paga (SPP), a government subsidy in Colombia that provided tuition support for post-secondary students to attend a government-certified higher education institution. Eligibility was based on both merit and economic need: in order to qualify for the program, students had to obtain a high grade in Colombia’s national standardized high school exit exam, SABER 11, and they also had to come from economically disadvantaged families, as measured by the survey-based wealth index SISBEN. The resulting BD design looks exactly as in Figure 1(b) with $\mathbf{X}_{i}=(X_{1i},X_{2i})=(\text{{SABER11}}_{i},\text{{SISBEN}}_{i})$. The authors analyzed it by considering each cutoff separately and then pooling all observations in the other dimension, leading to two univariate RD designs: one with score $\textit{SISBEN}_{i}$ in the sample $\text{{SABER11}}_{i}\geq 0$, and the other with score $\text{{SABER11}}_{i}$ in the sample $\textit{SISBEN}_{i}\geq 0$. This approach is equivalent to splitting the analysis along each of the two linear segments of $\mathcal{B}$ and then employing a pooling approach. This idea leads to the following local regression specification:

### v2

@@ 2.1 BD Designs in Practice
We searched the academic literature in multiple social sciences to learn how applied researchers implemented the BD design. We conducted our search in four general queries. In our initial Economics query, we included ten leading journals in Economics—Journal of Political Economy, Review of Economic Studies, Econometrica, American Economic Review, Quarterly Journal of Economics, Review of Economics and Statistics, American Economic Journal: Applied Economics, American Economic Journal: Economic Policy, American Economic Journal: Macroeconomics, and American Economic Journal: Microeconomics. In our initial Political Science query, we searched six leading Political Science journals—American Political Science Review, American Journal of Political Science, Journal of Politics, Political Science Research and Methods, British Journal of Political Science, and Quarterly Journal of Political Science. In both cases, we searched for terms like “boundary discontinuity”, “multivariate regression discontinuity”, and “multiscore regression discontinuity”, and focused our attention on papers published since $2015$. We also expanded our search to include other articles that were repeatedly cited by the articles collected in our initial queries.
@@
We also searched in the field of Education, as methodological researchers in that area were among the earliest scholars who discussed the multi-score RD design in non-geographic settings. For this search, we started with the seminal methodological papers by Reardon and Robinson (2012) and Wong et al. (2013), and collected papers that cited these papers or were cited by them. Finally, we also included some papers in environmental science to illustrate uses of the BD design at the intersection between the natural and social sciences. Our review is not meant to provide an exhaustive list of the empirical BD literature. This literature is vast and cannot be fully captured by our approach since we excluded several journals and restricted the time period. Our goal is simply to provide an overview of recent empirical work using the BD design across the social sciences to gauge the most common strategies used for empirical analysis.
@@
The results of our search are summarized in Table 5. Our search yielded $82$ published papers. Of these, we found that the overwhelming majority ($76$ articles or approximately 93%) used a pooling approach, focusing on the average effect along the entire boundary and using the univariate closest distance to the boundary as the running variable. Fewer than ten articles were explicit in accounting for the heterogeneity along the boundary; of these, only five reported effects for different points or segments along the boundary (Gonzalez, 2021; Grout et al., 2011; Keele and Titiunik, 2015; Snider and Williams, 2015; Velez and Newman, 2019). In sum, our review of the practical literature indicates that most applied researchers adopt a pooling approach, performing localization by grouping all observations sufficiently close to the boundary and then estimating a single average treatment effect. These common practices motivate the taxonomy we use in the rest of this chapter.
@@ 3 Pooling-Based Methods
Given a sample of outcomes and bivariate scores, $(Y_{1},\mathbf{X}_{1}),\ldots,(Y_{n},\mathbf{X}_{n})$, a distance function $\mathcal{d}(\cdot,\cdot)$ is used to measure the closeness between any two points in $\mathcal{X}$. The most common example is the Euclidean distance $\mathcal{d}(\mathbf{x}_{1},\mathbf{x}_{2})=\|\mathbf{x}_{1}-\mathbf{x}_{2}\|=\sqrt{(x_{11}-x_{21})^{2}+(x_{12}-x_{22})^{2}}$ for $\mathbf{x}_{j}=(x_{j1},x_{j2})$, $j=1,2$, although in some applications this is not the most appropriate measure. For example, Ambrus et al. (2020) are interested in how close residences in a city are to each other in terms of how long it takes for a person to walk between them. This requires calculating a walking distance, which is different from the Euclidean distance that cannot take into account, for example, that the path must follow city blocks and cannot go through buildings. More broadly, in geographic settings it is common practice to employ specific geodetic distances (Banerjee, 2005).
Once a distance function $\mathcal{d}(\mathbf{X}_{i},\mathbf{x})$ has been chosen, researchers who employ the pooling approach define the closest signed distance to the boundary:
@@
for $i=1,\ldots,n$. For a unit $i$, $D_{i}$ measures how far $i$’s location $\mathbf{X}_{i}$ is to the boundary point that is closest to $\mathbf{X}_{i}$, regardless of where that boundary point is located. Following the standard RD design logic, but now expanded to the BD design, localization occurs simultaneously along the entire assignment boundary because the univariate distance $D_{i}$ to $\mathcal{B}$ is used to construct the region such that $|D_{i}|\leq h$, which covers the boundary $\mathcal{B}$, where $h$ denotes a bandwidth parameter. This region, called a “tubular neighborhood” in mathematics (Federer, 2014), is a natural generalization of the standard localization approach used in univariate RD designs, where the region is simply $|X_{i}-c|\leq h$ when the score $X_{i}$ is scalar and $c$ is the univariate cutoff.
Figure 2 illustrates the idea graphically. In both cases, the tubular neighborhood represents the area that includes the observations used in the pooling-based analysis. Figure 2(a) shows a large bandwidth $h_{1}$, while Figure 2(b) shows a smaller bandwidth $h_{2}$. All units contribute simultaneously, regardless of their specific location. In particular, two units $i$ and $j$ with the same signed distance to the boundary, $D_{i}=D_{j}=d$, will be assigned to the same group (treatment or control, depending on the sign of $d$) and will be $|d|$ units away from the boundary, but could be far from each other.
@@
Figure 3 presents the two resulting RD plots (Calonico et al., 2015) in the pooling approach, based on the outcome data $Y_{i}$ and the univariate distance score $D_{i}$ (the cutoff is $c=0$ by construction of $D_{i}$). These plots include a fourth-order global polynomial fit (solid line) for visualization purposes, but these global approximations are not recommended for estimation of treatment effects because they tend to exhibit poor performance due to Runge’s phenomenon. Instead, as anticipated in Figure 2, the idea is to localize around the cutoff (i.e., the one-dimensional boundary curve) to conduct estimation and inference.
@@
In the pooling approach, the analysis is conducted using all the observations with score $\mathbf{X}_{i}$ within the tubular neighborhood determined by the bandwidth $h$, that is, for all units with $|D_{i}|\leq h$. However, within this subsample, the specific treatment effect estimation approach can vary substantially. To describe and unify the most common approaches in the empirical literature illustrated in Table 5 as well as other approaches recommended from a methodological perspective, we define the following generic weighted least squares regression notation:
@@
where $Y_{i}$ is the outcome variable, $\mathbf{Z}_{i}$ is a (column) vector of independent variables, and $W_{i}$ are weights. A localized regression analysis can be implemented by setting $W_{i}=\mathds{1}(|D_{i}|\leq h)$ or, more generally, $W_{i}=K(D_{i}/h)$ for a kernel function $K(\cdot)$ determining the relative weight given to each observation. The bandwidth $h$ determines the degree of localization, as illustrated in Figure 2. When $\mathbf{Z}_{i}$ contains polynomial expansions of $D_{i}$, the resulting estimation approach is often called nonparametric local polynomial regression but, as we will see shortly, empirical researchers often employ other variables in their local fit.
@@
All empirical approaches naturally include $T_{i}$ in $\mathbf{Z}_{i}$ to estimate treatment effects; the majority also include other variables such as $D_{i}$, $\mathbf{X}_{i}$, boundary-segment fixed effects (formally defined below), as well as certain transformations and interactions thereof. For simplicity, we omit discussing the inclusion of predetermined covariates, but we note that researchers often further augment the basic local regression specifications with pretreatment variables such as census characteristics, terrain ruggedness and elevation, time fixed effects, among other features. For a discussion of the role of preintervention covariates in the standard RD design see Calonico et al. (2019) for efficiency gains and Calonico et al. (2025) for treatment effect heterogeneity.
The pooling approach for analyzing BD designs begins by choosing the localization bandwidth $h$ to determine the units with $|D_{i}|\leq h$ that will be used in the subsequent estimation. Until now, there were no formal methods for choosing the bandwidth $h$ in an objective and data-driven way, so researchers often considered a range of ad-hoc values informed by practice and substantive knowledge. For example, Black (1999) reported results in three different bandwidths, keeping observations within $0.35$, $0.25$, and $0.15$ miles of the nearest point on the boundary. As in standard univariate RD designs, different choices of $h$ trade off bias and variance in the estimation, with a larger $h$ leading to more bias but less variance, and a smaller $h$ having the opposite effect. Cattaneo et al. (2026) discusses formal results characterizing this tradeoff and provides formal guidance for choosing the bandwidth when implementing pooling BD methods. We discuss these methods in Section 3.1.
Once the bandwidth $h$ has been chosen, researchers pool all units with distance $D_{i}$ within $h$ and then compare those assigned to the treatment group ($D_{i}\geq 0$) to those assigned to the control group ($D_{i}<0$). Because the bandwidth localizes the analysis to only a small region on either side of the boundary $\mathcal{B}$, the empirical analysis compares barely treated units to barely control units which, as in the standard univariate RD design, is the basis for the causal interpretation of the comparison, as all confounders are assumed to vary smoothly at the boundary while the treatment assignment changes abruptly from zero to one.
The most basic analysis proceeds by assuming that the units within the shrinking tubular neighborhood of width $2h$ are as-if randomly assigned to treatment and control. This approach is akin to the “local randomization” framework in standard RD designs (Lee and Lemieux, 2010; Cattaneo et al., 2015, 2017). A natural treatment effect estimator is the difference-in-means of the outcome between treated and control units with distance to the boundary no larger than the bandwidth:
@@
Black (1999, Table III) reported results using specification (1), calling it a “rough, nonparametric estimate of the value of better schools” (p. 590). She also augmented this basic specification to include “segment fixed effects”, that is, binary variables indicating whether an observation is closest to one of several segments partitioning the assignment boundary. To describe this approach formally, we let $\mathcal{B}_{\ell}$, $\ell=1,\ldots,L$, be disjoint segments that partition the assignment boundary, that is, $\mathcal{B}=\sqcup_{1\leq\ell\leq L}\mathcal{B}_{\ell}$. Then, $S_{i}=\operatorname*{arg\,min}_{1\leq\ell\leq L}d(\mathbf{X}_{i},\mathcal{B}_{\ell})$ is the boundary-segment allocation random index for each unit $i=1,\ldots,n$ (when the $\operatorname*{arg\,min}$ contains more than one element, the smallest one is chosen). The vector of segment fixed effects for observation $i$ is $\boldsymbol{\iota}_{L}(S_{i})=(\mathds{1}(S_{i}=1),\cdots,\mathds{1}(S_{i}=L))$, which collects the $L$ boundary segment binary indicators determined by $i$’s location. The difference-in-means specification with segment fixed effects is
@@
where the intercept is removed to prevent perfect collinearity. Setting $L=1$ reduces $\boldsymbol{\iota}_{L}(S_{i})$ to the common intercept in specification (1).
@@
Black (1999, Table II) used a modified version of specification (2) that included the test scores of the school associated with observation $i$ instead of the binary treatment assignment indicator $T_{i}$. The causal interpretation of this specification rests on the assumption that, after localizing the analysis to observations near the boundary of school attendance districts, the only factor that can explain a difference in average house prices between the treated and control areas is school quality, because school quality changes discontinuously at the boundary while all other determinants of house prices are equal on average. In this sense, using test scores instead of the treatment assignment indicator in specifications (1) and (2) is in the spirit of an instrumental variables strategy where $T_{i}$ is used as an instrument for test scores (which proxy school quality), and the localization to a small bandwidth around the boundary ensures that units assigned to treatment and units assigned to control are similar in all other characteristics that might affect house prices.
@@
The seminal research design used by Black (1999) is prior to the development of modern methods for RD analysis, which occurred predominantly during the $2000$s; see Cattaneo and Titiunik (2022) for an overview. In more recent practice, the simple specifications (1) and (2) are not commonly used because researchers tend to adopt a continuity-based approach rather than a local randomization approach for interpretation. Under the continuity-based approach, the local constant fit has poor control of the misspecification bias, and hence practitioners prefer more flexible local regression specifications.
@@
Later empirical work using BD designs made more explicit the connection between geographic discontinuities in treatment assignment and the classical RD design, following more closely the methods for univariate continuity-based RD analysis, which emphasized including the RD score in the polynomial specification to allow for arbitrary dependence between score and outcome even within the bandwidth. Although there is considerable diversity in the specifications used in more recent practice, most researchers who include the score adopted one of two main strategies: a polynomial fit of the outcome on the bivariate score $\mathbf{X}_{i}$, or a polynomial fit of the outcome on the univariate distance $D_{i}$.
@@
In an influential empirical study, Dell (2010) augmented specification (2) by including the bivariate score $\mathbf{X}_{i}$. She studied the long-run impact of the mita, a forced mining labor system instituted by the Spanish in Peru and Bolivia between $1573$ and $1812$. Dell used the historical boundary that determined which communities were forced to send labor to define treated and control areas, and studied the effect of the mita on contemporary outcomes measured at the household level, including consumption and childhood stunting. Her main local regression specification was
@@
which includes the treatment indicator, boundary segment fixed effects, and a polynomial expansion of the bivariate score $\mathbf{X}_{i}$ (with $p=3$ in her case). The estimated coefficient of interest is always the coefficient that accompanies the treatment indicator $T_{i}$. In our notation, $\mathbf{r}_{p}(\mathbf{u})=(u_{1},u_{2},u_{1}^{2},u_{1}u_{2},u_{2}^{2},\ldots,u_{1}^{p},u_{1}^{p-1}u_{2},\ldots,u_{1}u_{2}^{p-1},u_{2}^{p})$ with $\mathbf{u}=(u_{1},u_{2})$. The inclusion of (a polynomial basis of) the bivariate location score $\mathbf{X}_{i}$ is meant to control for “the smooth effects of geographic location” Dell (2010, p. 1875). A notable characteristic of this specification is that it does not include interactions between the bivariate score and the treatment indicator, which imposes the restriction that the polynomials be identical in the treated and control areas. For bandwidth, Dell used $100$ km, $75$ km and $50$ km within the mita boundary.
@@
Specification (3) has become a common strategy to analyze BD designs. Méndez and Van Patten (2022) use it to study the effect of private sector companies in the development of local amenities, using a land concession to the multinational United Fruit Company in Costa Rica. Paulsen et al. (2023) use it to estimate the effect of school funding on future earnings and political participation in New York, comparing areas that received higher funding to neighboring areas that did not. De Kadt and Larreguy (2018) compare areas in South Africa where traditional leaders remain highly influential to areas where they are not, and estimate the electoral impact of strong political ties between traditional leaders and the African National Congress (ANC). Other examples are listed in Table 5. All of these examples follow specification (2) by including a flexible polynomial expansion of latitude and longitude, as well as boundary-segment fixed effects, but without allowing for interactions with the treatment indicator $T_{i}$.
@@
with the polynomial expansion reducing to the simpler vector of the univariate distance $\mathbf{r}_{p}(D_{i})=(D_{i},D_{i}^{2},\ldots,D_{i}^{p})$. As in specification (3), specification (4) can also be augmented with boundary-segment fixed effects:
@@
where the estimated coefficient of interest continues to be the one accompanying $T_{i}$. In practice, specification (6) is often implemented with $p=1$ and, depending on whether standard software for RD design analysis is used, may employ a non-uniform weighting kernel (e.g., a triangular kernel that down-weights observations as their distance from the boundary increases). Putting aside the inclusion of boundary-segment fixed effects, Figure 6 illustrates the importance of including the interaction term $T_{i}\cdot\mathbf{r}_{p}(D_{i})$. Comparing Figures 5 and 6 shows that the two approaches can lead to different local approximations, with specifications (1) through (5) potentially exhibiting a larger bias than specification (6).
@@
Specification (6) is one of the most commonly used in recent practice. Examples include Ayres et al. (2021), Eugster et al. (2017), Jones et al. (2022a), Lalive (2008), and Lowes and Montero (2021). We note that several articles describe their empirical specifications as (4) or (5), but implement specification (6) in their analyses (as clarified in text, footnotes or replication files). These researchers employ standard practices and software from the univariate RD design literature, which follows specification (6).
@@
where the estimated treatment effect coefficient accompanying $T_{i}$ enjoys better bias properties due to the added flexibility in the local regression specification. However, this more flexible approach is not as common in the empirical literature—some exceptions are Gonzalez (2021), Egger and Wamser (2015), and Jones et al. (2022b).
In other applications, researchers estimate treatment effects in BD designs by considering two separate univariate RD designs, that is, keeping one dimension fixed and studying the other dimension of the bivariate score. For example, Londoño-Vélez et al. (2020) investigate the effects of Ser Pilo Paga (SPP), a government subsidy in Colombia that provided tuition support for post-secondary students to attend a government-certified higher education institution. Eligibility was based on both merit and economic need: in order to qualify for the program, students had to obtain a high grade in Colombia’s national standardized high school exit exam, SABER 11, and they also had to come from economically disadvantaged families, as measured by the survey-based wealth index SISBEN. The resulting BD design is as illustrated in Figure 1(b) with $\mathbf{X}_{i}=(X_{1i},X_{2i})=(\text{{SABER11}}_{i},\text{{SISBEN}}_{i})$. The authors analyzed it by considering each cutoff separately and then pooling all observations in the other dimension, leading to two univariate RD designs: one with score $\textit{SISBEN}_{i}$ in the sample $\text{{SABER11}}_{i}\geq 0$, and the other with score $\text{{SABER11}}_{i}$ in the sample $\textit{SISBEN}_{i}\geq 0$. This approach is equivalent to splitting the boundary along each of its two linear segments and then employing a pooling approach within each subsample. This idea leads to the following local regression specification: